find the slope and y-intercept of -x-4y=32.
Answer by checkley71(8403)

-X-4Y=32
-4Y=X+32
Y=X/-4+32/-4
Y=-X/4-8
SLOPE=-1/4.
Y INTERCEPT=-8.